OpenStudy (anonymous):

on my test last night i missed only questions where its the diffrence of 2 squares can someone explain to me what that means

OpenStudy (cwrw238):

an example is x^2 - 9 x^2 and 9 are 2 perfect squares to factor this you take the square root of each and plug these into standard form which is a^2 - b^2 = (a - b)( a + b)

OpenStudy (cwrw238):

so to factor x^2 - 9 - sq roots are x and 3: = (x - 3)( x + 3)